### v1

@@ 2.4 Alignment accuracy
Table 4: GraphAligner alignment results for PacBio HiFi and ONT reads aligned on LCPan and VG graphs from chromosomes 1, 10, and $2$.
@@ 4 Methods
Below, we describe a simple variant of LCP, expressed in terms of four rules, to identify cores in an input string from a given alphabet (e.g., the $4$-letter DNA alphabet). We will later show how to apply LCP iteratively using DCT in Section 4.1.
@@
Figure 2 depicts an example of the cores identified in a short DNA sequence. As the cores are identified, the two-bit encoding of the underlying characters (i.e., A = 00, C = 01, G = 10, T = 11) is used to form the core’s bitstream. Note that we use the term core alphabet to refer to the bitstreams that form the cores, specifically the DNA encoding used in the first LCP iteration.
@@
Adjacency Property: within a substring of length $3$, at most $2$ characters can be the starting positions of the cores.
@@ 4.1 Iterative Application of Locally Consistent Parsing
Our goal is to apply LCP iteratively to increase the distances between consecutive cores while reducing their number and increasing their length by a constant factor. However, since the core bitstreams are larger than the string’s alphabet (in our context, DNA), directly applying the LCP rules to these bitstreams would increase the distances without constraints. To address this, starting from the second LCP iteration ($i>1$), the bitstreams and LCP operations depend on the reduction in alphabet size via DCT. [^6] When fully applied, the DCT reduces the alphabet size to $2$, enabling a uniform bitstream length of $6$, which is crucial for maintaining consistency in later LCP iterations. After this reduction in alphabet size, the string of core bitstreams will satisfy both Lemmas 1 and 2, maintaining an even distance distribution and compact length; thus, LCP can be applied iteratively.
@@
Figure 3: DCT for reducing bitstreams to a new alphabet. Each block above corresponds to the bitstream of a core. The DCT compares each core’s bitstream to its left neighbor to form a shorter alphabet. E.g., the least significant bit of the core bitstream $11101011$, which differs from its left neighbor $011011$ at the fourth index (counting from the right and starting at index 0). The value of this bit is 0. Therefore, DCT replaces the core bitstream $11101011$ with the concatenation of the bits of $4$, and the value of the differing bit, resulting in $1000$. The figure shows the reduced alphabet inferred by the DCT for each core bitstream, on which LCP can be applied.
@@ Acknowledgments
Adjacency Property: within a substring of length $3$, there could be at most $2$ characters that could be the starting positions of cores.

### v2

@@ 2.4 Alignment accuracy
Table 4: GraphAligner alignment results for PacBio HiFi and ONT reads aligned on LCPan and VG graphs from chromosomes 1, 10, and 22.
@@ 4 Methods
Below, we describe a simple variant of LCP, expressed in terms of four rules, to identify cores in an input string from a given alphabet (e.g., the 44-letter DNA alphabet). We will later show how to apply LCP iteratively using DCT in Section 4.1.
@@
Figure 2 depicts an example of the cores identified in a short DNA sequence. As the cores are identified, the two-bit encoding of the underlying characters (i.e., A = $0$, C = 01, G = 10, T = 11) is used to form the core’s bitstream. Note that we use the term core alphabet to refer to the bitstreams that form the cores, specifically the DNA encoding used in the first LCP iteration.
@@
Adjacency Property: within a substring of length 33, at most 22 characters can be the starting positions of the cores.
@@ 4.1 Iterative Application of Locally Consistent Parsing
Our goal is to apply LCP iteratively to increase the distances between consecutive cores while reducing their number and increasing their length by a constant factor. However, since the core bitstreams are larger than the string’s alphabet (in our context, DNA), directly applying the LCP rules to these bitstreams would increase the distances without constraints. To address this, starting from the second LCP iteration ($i>1$), the bitstreams and LCP operations depend on the reduction in alphabet size via DCT. [^6] When fully applied, the DCT reduces the alphabet size to 22, enabling a uniform bitstream length of 66, which is crucial for maintaining consistency in later LCP iterations. After this reduction in alphabet size, the string of core bitstreams will satisfy both Lemmas 1 and 2, maintaining an even distance distribution and compact length; thus, LCP can be applied iteratively.
@@
Figure 3: DCT for reducing bitstreams to a new alphabet. Each block above corresponds to the bitstream of a core. The DCT compares each core’s bitstream to its left neighbor to form a shorter alphabet. E.g., the least significant bit of the core bitstream $11101011$, which differs from its left neighbor $011011$ at the fourth index (counting from the right and starting at index 0). The value of this bit is 0. Therefore, DCT replaces the core bitstream $11101011$ with the concatenation of the bits of 44, and the value of the differing bit, resulting in $1000$. The figure shows the reduced alphabet inferred by the DCT for each core bitstream, on which LCP can be applied.
@@ Acknowledgments
Adjacency Property: within a substring of length 33, there could be at most 22 characters that could be the starting positions of cores.